# Two new universal inequalities for Neumann eigenvalues of the Laplacian on a planar convex domain

### v2

@@ 1. Introduction
where $\partial_{n}$ denotes the directional derivative with respect to $n$, the outward unit normal vector to $\partial\Omega$. In this situation it is known the compactness of the embedding $H^{1}(\Omega)$ to $L^{2}(\Omega)$ and it gives that the eigenvalues of (1.1) can be listed with multiplicity as follows:
@@
We also obtain the opposite inequality:
@@
In the proof of the above main theorems we use a variant of domain monotonicity for eigenvalues to reduce the problem in the case of rectangular domains. In this aspect we give a new brief proof for an weaker inequality of the inequality (1.2) in the last section (see Theorem 4.1). It is weaker but its proof is new and rather shorter than the proof of the inequality (1.2) so the author include it in this paper.
@@ 2. Preliminaries
Let $\Omega$ be a bounded convex domain in $\mathbb{R}^{n}$. Tlanslating $\Omega$ if necessary we can find a sequence $a_{1},a_{2},\cdots,a_{n}$ of positive real numbers such that
@@
Combining the above proposition with the next theorem we see that the proof of the main theorem reduce to the case where the domain is rectanglular. It is a variant of domain monotonicity, it states that the smaller domain has a larger eigenvalue up to a multiplicative constant factor under convexity assumption.
@@
For a convex domain Neumann eigenvalues of the Laplacian are related with diameter as follows:
@@
Theorem 2.6 (Buser, [1, 8.2.1 Theorem]).
@@ 3. Proof of the Main Theorems
Put $R:=\frac{Ck}{l\sqrt{\mu_{k}(\Omega)}}$, where $C>0$ is a sufficiently large constant which will be specified later. We want to show that $\Omega$ can be decomposed into $l^{\prime}$ convex subdomains $\{\Omega_{i}\}_{i=1}^{l^{\prime}}$ such that $\operatorname{Diam}\Omega_{i}^{\prime}\lesssim R$ and $l^{\prime}\leq l$. If this holds, then using Buser’s theorem (Theorem 2.6) and the Payne-Weinberger inequality (Theorem 2.3) we obtain
@@
We first consider the case where $a\leq 2R$. In this case since Kröger’s inequality (Theorem 2.4) gives
$R\gtrsim\frac{Ck}{l\frac{k}{\operatorname{Diam}\Omega}}=\frac{C\operatorname{Diam}\Omega}{l}\gtrsim\frac{C(2b)}{l},$
if $C>0$ is large enough, then we can have a partition
@@
such that $l^{\prime}\leq l$ and $b_{i+1}-b_{i}\leq R$. Thus setting $\Omega_{i}:=[-a,a]\times[b_{i},b_{i+1}]$ we get the desired convex partition of $\Omega$; $\Omega=\bigcup_{i=1}^{l^{\prime}}\Omega_{i}$, where $\operatorname{Diam}\Omega_{i}\lesssim R$.
@@
Note that $\Omega^{\prime}$ is a nonempty (noncollapsed) rectangular region. Let $\{x_{i}\}_{i=1}^{l^{\prime}}$ be a maximal $(2R)$-separated points in $\Omega^{\prime}$. Since $\{B(x_{i},R)\}_{i=1}^{l^{\prime}}$ are disjoint balls in $\Omega$, we have
@@
If $C$ is large enough we then obtain $l^{\prime}\leq l$.
@@
then $\{\Omega_{i}\}_{i=1}^{l^{\prime}}$ is a convex partition of $\Omega$. Since $\Omega_{i}\subseteq B(x_{i},(2+\sqrt{2})R)$ we obtain the desired convex partition $\{\Omega_{i}\}_{i=1}^{l^{\prime}}$ of $\Omega$. This completes the proof of the theorem. $\Box$
@@
Put $R:=\frac{C\sqrt{l}}{\sqrt{k\mu_{l}(\Omega)}}$, where $C>0$ is a large constant which will be specified later. With the same reason in the proof of Theorem 1.1 it suffices to prove the existence of a convex partition $\{\Omega_{i}\}_{i=1}^{k^{\prime}}$ of $\Omega$ such that $k^{\prime}\leq k$ and $\operatorname{Diam}\Omega_{i}\lesssim R$ for each $i$.
@@
If we take large enough $C$ then we get $R\geq\frac{2b}{k}$ and the same argument in the Theorem 1.1 shows the existence of the desired convex partition.
@@
and take a maximal $2R$-separated points $\{x_{i}\}_{i=1}^{k^{\prime}}$ in $\Omega^{\prime}$. If we show $k^{\prime}\leq k$ then as in the proof of Theorem 1.1 we get the desired convex partition $\{\Omega_{i}\}_{i=1}^{k^{\prime}}$. Since $\{B(x_{i},R)\}_{i=1}^{k^{\prime}}$ are disjoint balls in $\Omega$, we have
@@
For sufficiently large $C$ we thereby obtain $k^{\prime}\leq k$. This completes the proof. $\Box$
@@
Proposition 2.1 gives the existence of a rectangular domain $R$ such that
@@
[…]
$\mu_{k+1}(R)\lesssim\mu_{k}(R).$
Applying Theorem 2.2 twice we thus obtain
@@
This completes the proof. $\Box$
Acknowledgement. The author would like to thank Arataka Funano and Chizu Funano for their massive support. This work was supported by JSPS KAKENHI Grant Number JP24K06731.

### v3

@@ Abstract.
Résumé. Nous établissons deux nouvelles inégalités universelles pour les valeurs propres de Neumann du laplacien sur un domaine convexe plan.
@@ 1. Introduction
where $\partial_{n}$ denotes the directional derivative with respect to $n$, the outward unit normal vector to $\partial\Omega$. In this situation the compactness of the embedding $H^{1}(\Omega)\subset L^{2}(\Omega)$ implies that the eigenvalues of (1.1) can be listed with multiplicity as follows:
@@
We also obtain a reverse inequality:
@@
In the proof of the above main theorems we use a variant of domain monotonicity for eigenvalues to reduce the problem to the case of rectangular domains. In this aspect we give a new brief proof for a weaker version of the inequality (1.2) in the last section (see Theorem 4.1). It is weaker but its proof is new and rather shorter than the proof of the inequality (1.2) so the author includes it in this paper.
@@ 2. Preliminaries
Let $\Omega$ be a bounded convex domain in $\mathbb{R}^{n}$. Choosing the coordinate axes appropriately if necessary we can find a sequence $a_{1},a_{2},\cdots,a_{n}$ of positive real numbers such that
@@
Combining the above proposition with the next theorem we see that the proof of the main theorem reduces to the case where the domain is rectanglular. It is a variant of domain monotonicity, it states that the smaller domain has a larger eigenvalue up to a multiplicative constant factor under convexity assumption.
@@
For a convex domain, Neumann eigenvalues of the Laplacian are related with the diameter as follows:
@@
Theorem 2.6 (Buser, [1, Theorem 8.2.1]).
@@ 3. Proof of the Main Theorems
Put $R:=\frac{Ck}{l\sqrt{\mu_{k}(\Omega)}}$, where $C>0$ is a sufficiently large constant which will be specified later. We want to show that $\Omega$ can be decomposed into $l^{\prime}$ convex subdomains $\{\Omega_{i}\}_{i=1}^{l^{\prime}}$ such that $\operatorname{Diam}(\Omega_{i})\lesssim R$ and $l^{\prime}\leq l$. If this holds, then using Buser’s theorem (Theorem 2.6) and the Payne-Weinberger inequality (Theorem 2.3) we obtain
@@
We first consider the case where $a\leq 2R$. In this case Kröger’s inequality (Theorem 2.4) gives
$R\gtrsim\frac{Ck}{l\frac{k}{\operatorname{Diam}\Omega}}=\frac{C\operatorname{Diam}\Omega}{l}\geq\frac{Cb}{l}.$
We can take an absolute constant $C_{1}$ such that if $C\geq C_{1}$, then we have $R\geq 2b/l$. We thereby have a partition
@@
such that $l^{\prime}\leq l$ and $b_{i+1}-b_{i}\leq R$. Thus setting $\Omega_{i}:=[-a,a]\times[b_{i},b_{i+1}]$ we get the desired convex partition of $\Omega$; $\Omega=\bigcup_{i=1}^{l^{\prime}}\Omega_{i}$, where $\operatorname{Diam}(\Omega_{i})\lesssim R$.
@@
Note that $\Omega^{\prime}$ is a nonempty (noncollapsed) rectangular region. Let $\{x_{i}\}_{i=1}^{l^{\prime}}$ be a maximal set of $2R$-separated points in $\Omega^{\prime}$. Since $\{B(x_{i},R)\}_{i=1}^{l^{\prime}}$ are disjoint balls in $\Omega$, we have
@@
There exists an absolute constant $C_{2}>0$ such that if $C\geq C_{2}$ we then obtain $l^{\prime}\leq l$.
@@
then $\{\Omega_{i}\}_{i=1}^{l^{\prime}}$ is a convex partition of $\Omega$. Since $\Omega_{i}\subseteq B(x_{i},(2+\sqrt{2})R)$ we obtain the desired convex partition $\{\Omega_{i}\}_{i=1}^{l^{\prime}}$ of $\Omega$.
Taking $C$ as the maximum of $C_{1}$ and $C_{2}$ completes the proof of the theorem. $\Box$
@@
Put $R:=\frac{C\sqrt{l}}{\sqrt{k\mu_{l}(\Omega)}}$, where $C>0$ is a large constant which will be specified later. Using the same reasoning as in the proof of Theorem 1.1 it suffices to prove the existence of a convex partition $\{\Omega_{i}\}_{i=1}^{k^{\prime}}$ of $\Omega$ such that $k^{\prime}\leq k$ and $\operatorname{Diam}(\Omega_{i})\lesssim R$ for each $i$.
@@
There exists an absolute constant $C_{1}>0$ such that if $C\geq C_{1}$ then we get $R\geq\frac{2b}{k}$. Therefore the same argument as in the Theorem 1.1 shows the existence of the desired convex partition.
@@
and take a maximal set of $2R$-separated points $\{x_{i}\}_{i=1}^{k^{\prime}}$ in $\Omega^{\prime}$. If we show $k^{\prime}\leq k$ then as in the proof of Theorem 1.1 we get the desired convex partition $\{\Omega_{i}\}_{i=1}^{k^{\prime}}$. Since $\{B(x_{i},R)\}_{i=1}^{k^{\prime}}$ are disjoint balls in $\Omega$, we have
@@
We can take an absolute constant $C_{2}>0$ such that if $C\geq C_{2}$ we then obtain $k^{\prime}\leq k$.
Setting $C:=\max\{C_{1},C_{2}\}$ completes the proof. $\Box$
@@
Proposition 2.1 gives the existence of a rectangular domain $R=[-a_{1},a_{1}]\times[-a_{2},a_{2}]\times\cdots\times[-a_{n},a_{n}]$ such that
@@
We claim that $\mu_{k+1}(R)\lesssim\mu_{k}(R)$. If this claim holds, applying Theorem 2.2 twice, we then obtain
@@
which implies the conclusion of the theorem.
To prove the claim, let $\widetilde{R}$ be the flat torus whose fundamental domain is $R$. Recall that Neumann eigenvalues of the Laplacian on $R$ are of the form
$\frac{\pi^{2}}{4}\Big\{\frac{k_{1}^{2}}{a_{1}^{2}}+\frac{k_{2}^{2}}{a_{2}^{2}}+\cdots+\frac{k_{n}^{2}}{a_{n}^{2}}\Big\},\ k_{1},k_{2},\cdots,k_{n}\in\mathbb{N}\cup\{0\}$
[…]
$4\pi^{2}\Big\{\frac{l_{1}^{2}}{a_{1}^{2}}+\frac{l_{2}^{2}}{a_{2}^{2}}+\cdots+\frac{l_{n}^{2}}{a_{n}^{2}}\Big\},\ l_{1},l_{2},\cdots,l_{n}\in\mathbb{Z}$
([2, Section II.2]). Suppose that
$\mu_{k}(R)=\frac{\pi^{2}}{4}\Big\{\frac{m_{1}^{2}}{a_{1}^{2}}+\cdots+\frac{m_{n}^{2}}{a_{n}^{2}}\Big\}\text{ and }\mu_{k+1}(R)=\frac{\pi^{2}}{4}\Big\{\frac{m_{1}^{\prime 2}}{a_{1}^{2}}+\cdots+\frac{m_{n}^{\prime 2}}{a_{n}^{2}}\Big\}$
for some $m_{1},\cdots,m_{n},m_{1}^{\prime},\cdots,m_{n}^{\prime}\in\mathbb{N}\cup\{0\}$. Then we can take $l$ such that
$\lambda_{l}(\widetilde{R})=4\pi^{2}\Big\{\frac{m_{1}^{2}}{a_{1}^{2}}+\cdots+\frac{m_{n}^{2}}{a_{n}^{2}}\Big\}\text{ and }\lambda_{l+1}(\widetilde{R})=4\pi^{2}\Big\{\frac{m_{1}^{\prime 2}}{a_{1}^{2}}+\cdots+\frac{m_{n}^{\prime 2}}{a_{n}^{2}}\Big\}.$
[…]
Acknowledgement. The author would like to thank Arataka Funano and Chizu Funano for their massive support. The author also thanks the anonymous referee for his or her suggestion. This work was supported by JSPS KAKENHI Grant Number JP24K06731.